Simplify the following expression 42a4 b¹²/ 7a² b³, then fill in the values for A, B and C as indicated


Simplify the following expression (5x² + x(5 - 4x) + 1-5x + x2 - 9)/ 24x, then fill in the values of A, B and C as indicated


Simplify the following expression 2x(4- 2x + y) - 5x(y² + x - y), then fill in the answers for A, B, C and D as indicated


1
Expert's answer
2021-06-22T05:58:20-0400

(A)

"\frac {42a^4b^{12}}{7a^2b^3} \\

6a^{4-2}b^{12-3} \\

6a^2b^9 \\"

(B)

"\frac {(5x^2+x(5-4x)+1-5x+x^2-9)}{24x} \\

\frac {(6x^2+5x-4x^2+1-5x-9)}{24x} \\

\frac {2x^2-8}{24x} \\

\frac {x^2-4}{12x} \\

\frac {(x+2)(x-2)}{12x} \\"

(C)

"2x(4-2x + y) - 5x(y^2 + x - y) \\

8x-4x^2+2xy-5xy^2-5x^2+5xy \\

8x-9x^2+7xy-5xy^2"